If f(x) = 2x +5 and g(x) = x – 3, find the inverse of f(g(x)) and write your
answer as a single fraction in reduced form.

Answer by stanbon(75887)

If f(x) = 2x +5 and g(x) = x – 3, find the inverse of f(g(x)) and write your
answer as a single fraction in reduced form.
---------------
[f(g(x)]^-1 = g^-1(f^-1(x))
----------------------------
f^-1 = (x-5)/2
g^-1 = x+3
--------------------
So [f(g(x)]^-1 = g^-1[(x-5)/2] = [(x-5)/2]+3 = [x-5+6]/2 = [x+1]/2
======================================================================
Cheers,
Stan H.